OpenStudy (anonymous):

What is the range of the graph of y = 3(x – 3)^2 + 4?

zepdrix (zepdrix):

This function represents a Parabola in standard form. A parabola, OPENING UPWARD, looks like this,\[\huge y=x^2\]Our function has been stretched by a factor of 3,\[\huge y=3x^2\]It has had it's vertex shifted to the RIGHT 3 units,\[\huge y=3(x-3)^2\]And the vertex has also been shifted UP 4 units,\[\huge y=3(x-3)^2+4\]So that's where all of those numbers are coming from. Remember what a parabola looks like? It's like a bowl shape, in this case it's opening upward with it's lowest point at (3,4).

zepdrix (zepdrix):

The range represents the possible Y values. So if 4 is our LOWEST possible Y value, and it takes on every Y value larger than that, how could we write that in interval notaion? or whatever notation you're using in class :)

OpenStudy (anonymous):

y less than or equal to 4 y greater than or equal to 4 y less than or equal to 3 y greater than or equal to 3

zepdrix (zepdrix):

The lowest point of this parabola is at (3,4). So it's defined for all values larger than the Y value of that coordinate pair.

zepdrix (zepdrix):

Lol, if you can't figure it out from all that info, you need to do some more homework... :CC

OpenStudy (anonymous):

Thanks I Understand now had to reread :D